OpenStudy (anonymous):

can someone explain the difference between du/dv and dv/dx so I can understand the chain rule better??

OpenStudy (campbell_st):

ok.... if you have y = (2x + 5)^5 the way to approch it is to do some substitution let u = 2x + 5 so you are looking at y = u^5 find the derivative of y with respect to u dy/du dy/du = 5u^4 but you also need to deal with the 2x + 5 since it is equal to u so you need to differentiate u with respect to x du/dx du/dx = 2 so to find dy/dx you need to use the chain rule dy/dx = dy/du * du/dx = 5(u)^4 * 2 make the substitution for u dy/dx = 10(2x + 5)^4 hope this helps

OpenStudy (anonymous):

Thank you so much!